A GENERAL CLASS OF CHANGE POINT AND CHANGE CURVE MODELING FOR LIFE TIME DATA

KAUSHIK PATRA^{1} AND DIPAK K. DEY^{2}

1Center for Biostatistics in AIDS Research, Department of Biostatistics,

Harvard School of Public Health, 5th Floor, 651 Huntington Ave, Boston, MA 02115, U.S.A.

2Department of Statistics, University of Connecticut, U120, 215 Glen Brook Road, Storrs, CT 06269, U.S.A.

(Received August 28, 2000; revised March 19, 2001)

Abstract. Change point hazard rate models arise in many life time data analysis, for example, in studying times until the undesirable side effects occur in clinical trials. In this paper we propose a general class of change point hazard model for survival data. This class includes and extends different types of change point models for survival data, e.g. cure rate model and lag model. Most classical approach develops estimates of model parameters, with particular interest in change point parameter and often the whole hazard function, but exclusively in terms of asymptotic properties. We propose a Bayesian approach, avoiding asymptotics and provide inference conditional upon the observed data. The proposed Bayesian models are fitted using Markov chain Monte Carlo method. We illustrate our proposed methodology with an application to modeling life times of the printed circuit board.

Key words and phrases: Change point, Gibbs sampling, hazard function, posterior inference, survival function.